Title

Article text.

INAF queue observing, 2016 December 3 - 4

Observer: R. Carini, A. Marchetti
Telescope Operator: G. Bechetti
Support Astronomer: D. Thompson
Instrument: LUCI1, MODS1, MODS2

Summary

At sunset the sky is clear, but humidity is at 95%, and temperature very near to dew point.

LUCI2 is not available

00:55 humidity 100%, temperature at the dew point.
the dome is closed
during the whole night the seeing was very bad, oscillating between 1.5'' e 2.5'', with peaks to 3''

Night duration (12\xB0 el twilight) 18:10-6:00 ( 01:10-13.0 UT) = 11h 50m

Observing time: 6h 15m (53%)

Weather time loss: 5h 35m (47%)

technical time loss 0m (0%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
LUCI 1 Antoniucci, Marziani B DARKS      
LUCI 1 Antoniucci, Marziani B FLATS      
LUCI 1 Antoniucci, Marziani B ARCS      
MODS 1-2 Pancino Pandas-17     seeing out of spec, misunderstanding between observers, the target has been already taken
MODS 1-2 Antoniucci V1118Ori dual grating 20 m spec COMPLETED
MODS 1-2 Antoniucci V1118Ori blue : g_sdss; red: r_sdss 2m imaging COMPLETED
MODS 1-2 Standard Feige34 dual grating    
LUCI1 Marziani J SDSS J1054+2536 G210HiRes H 0.5 h seeing out of spec, the frames taken in twilight from the third
LUCI1 SKy FLAT   Ks, H    

Details

LUCI1 darks for Antoniucci (8x2sec LIR)

luci1.20141204.0001-0005 camera 1.8

luci1.20141204.0006-0010 camera 3.75

LUCI1 darks for Marziani B (360sec MER)

luci1.20141204.0011-0015

LUCI1 Flat Field for Marziani B

luci1.20141204.0016-0017-0018- BAD (AGW was not homed)

luci1.20141204.0019- LOW COUNTS

luci1.20141204.0020-0024

LUCI1 Flat Field for Antoniucci

luci1.20141204.0025-0029 ZJ filter

luci1.20141204.0030-0031 HK filter, too high

luci1.20141204.0032-0036 HK filter, OK

Lamps for Antoniucci HK filter

luci1.20161204.0037 LOW COUNTS

luci1.20161204.0038-0040

Lamps for Antoniucci zJ filter

luci1.20161204.0041-0042

Lamps for Marziani B (G210 HiRes lambda 1.66 filter H)

luci1.20161204.0043-0044

06:30 The TO opens the dome

06:40 We wake up MODS1 and MODS2

06:50 The collimation had success

PANCINO PROGRAM

07:00 We start acquisition of Pancino Target PAndAs -17 in binocular mode

mods1r.20161204.0001-0003

mods2r.20161204.0001-0003

07:22 We start the science of Pandas-17

mods1r.20161204.0004-0005

mods1b.20161204.0001-0002

mods2r.20161204.0004-0005

mods2b.20161204.0001-0002

07:41 the seeing is about 1.6, we slew to Antoniucci back up program

ANTONIUCCI PROGRAM

07:49 acquisition of V1118Ori

the guide star is outside the patrol field after the telescope preset

08:16 after changing the parallactic angle in the script, we acquire V1118Ori again

mods1r.20161204.0006-0008

mods2r.20161204.0006-0008

08:33 we execute the science script:

mods1r.20161204.0009-0010

mods1b.20161204.0003-0004

mods2r.20161204.0009-0010

mods2b.20161204.0003-0004

08:55 we do v1118Ori imaging

mods1r.20161204.00011-0012

mods1b.20161204.0005-0006

mods2r.20161204.0011-0012

mods2b.20161204.0005-0006

we repeat with a lower exptime

mods1r.20161204.0013-0014

mods1b.20161204.0007-0008

mods2r.20161204.0013-0014

mods2b.20161204.0007-0008

09:26 we acquire the standard for MODS spectroscopy Feige34

==mods1r.20161204.0015-0016 aborted

09:41 we acquire again

mods1r.20161204.0017-0018

mods2r.20161204.0015-0016

09:54 we execute science script for Feige34

mods1r.20161204.0019-0020

mods1b.20161204.0009-0011 for some reason the MODS1 blue channel took one extra exposure

mods2r.20161204.0017-0018

mods2b.20161204.0009-0010

10:07 seeing has raised up to 2.5". There are also wind gusts against the secondary mirror

10:18 seeing is rapidly variable between 1.5" and 3"

10:42 we switch back to LUCI1

MARZIANI PROGRAM

11:28 we execute the telluric for Marziani J

luci.20161204.0045-0056 BAD * problem during the acquisition. In LUCI1 the LS0.75 is called LS_1.50arsec.In the script the PI the name of the slit is LS_0.75, so has benne created a conflict. the AO fixed t * the second problem during the acquisition of the telluric: in the script are not correct the filters.

luci.20161204.0057-0060 alignment

luci.20161204.0061-0063 Telluric (there is one more exposition because first is too near to the border, due to an accidental extra-offset)

12:27 we send the target script for Marziani J (SDSS J1054+2536)

luci.20161204.0064-0068 alignment

luci.20161204.0069-0073 Science. From 0071 on the frames are taken in twilight

The seeing was really bad, oscillating between 1.7 and 3.2.

Flat field of sky for LUCI (Ks, H2)

-- %USERSIG{AlidaMarchetti - 2016-12-04}%

Comments

 
Topic revision: r19 - 08 Dec 2016, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback